# RevCD - Reversed Conditional Diffusion for Generalized Zero-Shot Learning

### v1

@@ 3.4 Model Architecture
Our denoising Diffusion model employs a U-net architecture, as introduced by the probabilistic diffusion model in [19]. To merge visual and semantic information effectively, we have customized this architecture to support both our time-dependent and visual-dependent embeddings, as illustrated in Fig 2(b). To our knowledge, this represents the first application of a U-net architecture tailored for zero-shot learning in such a specific way. The encoder-decoder structure of our U-net is built from linear blocks featuring ReLU non-linearity and batch normalization. Inputs to each layer include sinusoidal time embeddings and conditional data, which are extracted using self-attention mechanisms and augmented by a skip-connection between the encoding and decoding stages.
@@
for $s_{0},\mathbf{x}\sim p(x,s)\in\mathcal{D}^{seen}$ do
@@
$\nabla_{\theta}\ [(1-\lambda)\mathcal{L}_{Diff}(\hat{s}_{0})+\lambda\mathcal{L}_{cls}(\hat{y})+\eta\mathcal{L}_{noise}(\hat{\epsilon}_{t})]$
3.5 Training Objective
Our main idea focuses on directly modeling the semantic posterior using variational inference rooted in Eq. (13). We achieve this by disentangling the posterior estimation into three key components: noise prediction, data reconstruction, and classification. This decomposition results in a more complex and nuanced loss landscape [29]. Despite the increased complexity, integrating these distinct loss components enhances the model’s generalization capabilities. This is primarily due to the regularization effects inherent in the multi-faceted loss function and the fine-tuning achieved through careful hyperparameter optimization.
$\begin{split}\mathbf{\mathcal{L}_{total}}=\\ (1-\lambda)\mathcal{L}_{diffusion}+\lambda\mathcal{L}_{classification}+\eta\mathcal{L}_{noise}\end{split}$ (13)
Here, $\lambda$ serves as a balancing factor between the objectives of reconstruction and classification, while the hyperparameter $\eta$ acts as a regularization coefficient. Through this, the probability distribution of the samples aligns with the expectation of the generated conditional samples $p(s)\propto\mathbb{E}_{x\sim p(x)}[p_{\theta}(s|x)$. The implementation details of this loss function during training are provided in Algorithm (1).
@@ 3.6 Sampling
Using standard methods from diffusion theory [19], we generate the semantic embedding space of an unseen sample through iterative conditional denoising using our trained model, as shown in Algorithm (2). Samples are drawn from the standard normal prior $p(s_{T})\sim\mathcal{N}(0,\mathbf{I})$ and denoised conditioned on the sinusoidal time-step embedding $\bar{t}_{i}\ \forall i\in[1000,0]$ and the Transformer-encoded latent visual space $x\in\mathcal{D}^{unseen}$.
@@
$s_{t}\sim\mathcal{N}(0,\mathbf{I})$
$x\sim p(x|y)\in\mathcal{D}^{unseen}$
@@
     $\hat{s}_{t}\leftarrow Unet(s_{t},x\oplus s,\bar{t})$
@@ 3.7 Zero-Shot Inference
In the zero-shot learning setting, the model utilizes the approximated semantic embeddings $\hat{s}^{unseen}$ to classify instances of unseen classes using a nearest-neighbor approach in the semantic space. Leveraging the semantic embeddings to bridge the gap between visual features of $x^{unseen}$ and class labels $y^{unseen}$:
$\hat{y}=\arg\min_{y\in\mathcal{Y}_{unseen}}\text{distance}(\hat{s}^{unseen},s^{unseen}_{y}),$ (15)
where $\hat{y}$ is the predicted class label for an unseen class instance, and $\text{distance}(\cdot,\cdot)$ is a distance metric, in our case cosine similarity:
@@ 4 Experimental results
Dataset. Our analysis of diffusion as a generative method for zero-shot inference employs four publicly available benchmark datasets, distinct within the field. This allows us to make a fair comparison of the quality and coverage approximated semantics. The datasets are: Scene Understanding Attribute dataset (SUN) [36], Caltech Birds dataset (CUB200-2011) [45], Animals with Attributes 2 dataset (AwA2) [52], and Attribute Pascal and Yahoo dataset (APY).
@@
The CUB dataset, focused on bird species, offers detailed representations in both image and semantic spaces. In contrast, AWA, which covers animal species, provides coarser descriptions in both domains. SUN, a scenery dataset, spans a wide range of classes, while APY consists of general objects with a limited, broad semantic description.
@@
Visual features are derived using a ResNet101 backbone pre-trained on ImageNet [50]. We only compare models using similar image features to ensure equitable. The semantic attributes released with the respective datasets are implemented, consisting of crowd-sourced human studies or word2vector label extractions.
@@
Implementation details. The employed U-net architecture for our Diffusion model consists of three hidden, fully connected dense layers, ReLU activation function, and dropout for regularization. We use a feature extractor with a multi-head self-attention layer (MSA) for the conditional space. In the encoder and the decoder of the U-net, we concatenate and add the sinusoidal time embedding to layer inputs and the conditional features as explained in section 3.1.1.
@@
Generative model Name AwA CUB SUN aPY Seen Unseen Harm. Seen Unseen Harm. Seen Unseen Harm. Seen Unseen Harm. VAE cVAE [7] 72.6 54.4 62.2 59.9 47.0 52.7 - - - 55.3 30.2 39.0 GAN GAN [13] 82.4 24.7 38.1 44.4 31.3 36.8 43.3 29.0 31.4 50.0 25.1 33.4 Diffusion (ours) RevCD 82.4 49.2 63.8 57.4 40.2 47.3 58.2 51.2 54.4 63.6 36.7 46.5
@@
4.1 Generalized accuracy
These comparisons are summarized in Table 1. As shown, no single generative model consistently outperforms the others across all datasets when generating both seen and unseen samples. Notably, our method surpasses the other approaches in generating samples on three of the four datasets when measured by the harmonic mean. The most significant performance gap is observed in the semantically coarse SUN dataset, where our approach achieves a 20% improvement over GANs. A similar trend is evident in the class-diverse APY and AWA datasets, albeit with smaller margins. In contrast, the CUB dataset, which features a wide variety of fine-grained semantic details, proves challenging for denoising approaches, making variational inference methods a more effective fit.
VAEs benefit from the tractable estimation of the posterior distribution, as evidenced by a 5.4% higher harmonic mean when both seen and unseen samples are drawn from tighter distributions, such as those observed in the CUB dataset, which emphasizes local descriptions. In contrast, GANs appear to struggle in this context, likely due to mode collapse in the posterior. Our Diffusion model performs moderately, achieving a 10.4% improvement over GANs.
Conversely, GANs implicitly learn the distribution through adversarial training, which encourages the generator to produce high-fidelity samples, as demonstrated in the AWA dataset, where attributes emphasize global image descriptions. However, the absence of explicit density estimation makes GANs susceptible to seen-unseen bias, leading to a preference for the seen distribution during inference, as observed in both the SUN and APY datasets. Our Diffusion model shows strong performance in generating samples when the seen distribution is discriminative and low-dimensional, as in the AWA and APY datasets. However, it struggles to maintain a tight lower bound on the true data distribution as the dimensionality of the semantic space increases, as evidenced by its performance in the CUB dataset.
Table 2 compares our Diffusion model with top-performing established zero-shot classification approaches. For SUN, our Diffusion model outperforms other approaches in the generation of the seen sample by a margin of 21.6%, with only a minor deficiency in unseen generation of 1.1% compared to f-CLSWGAN [51] and DAZLE [22] respectively. For CUB and APY, our model performs comparably to the other models. The pattern of the results indicates that generative models effectively outperform embedding-based approaches, with our Diffusion model being no exception.
Table 2: Comparative table for previously established methods in the field. The reported results are the generalized zero-shot per-class average results for seen $(S)$, unseen $(U)$ and harmonic mean $(H)$ as reported in the respective papers.
AWA2 CUB SUN APY Methods S U H S U H S U H S U H Embedding approach ALE [1] 81.8 14.0 23.9 62.8 23.7 34.4 33.1 21.8 26.3 - - - LATEM [49] 77.3 11.5 20.0 57.3 15.2 24.0 28.8 14.7 19.5 - - - PQZSL [30] - - - 43.2 51.4 46.9 35.1 35.3 35.2 64.1 27.9 38.8 Generative approach f-CLSWGAN [51] 68.9 52.1 59.4 57.7 43.7 49.7 36.6 42.6 39.4 61.7 32.9 42.9 CADA-VAE [39] 75.0 55.8 63.9 53.5 51.6 52.5 35.7 47.2 40.6 - - - DAZLE [22] 75.7 60.3 67.1 59.6 56.7 58.1 24.3 52.3 33.2 - - - RevCD (Ours) 82.4 49.2 63.8 57.4 40.2 47.3 58.2 51.2 54.4 63.6 36.7 43.1
Toy experiment: The effect of a lower dimension space. We explore how the implicit density modeling framework of our Diffusion model captures the distribution of the semantic space in a reduced setting. This ability allows our model to be implemented in environment where capturing broad auxiliary information is costly or infeasible. Fig. 4 exemplifies this by reducing the dimensions of the semantic space by capturing the principal components of the respective datasets and comparing them to the generative capabilities of a conditional VAE and a conditional GAN. We observe that our Diffusion model maintains solid performance in the generative estimation of unseen classes when only condensed information is available. Only a 2% decrease can be seen as we reduce the dimensions in CUB and 15% in SUN. We hypothesise that the cross Hadamard-Addition embedding captures available information during iterative denoising. This reduces the dependence on a single condition at any given time and allows the generative process to gradually be guided in the right direction.
Figure 4: We measure the effectiveness in terms of accuracy of unseen classification through nearest neighbour search of pure generative models VAE, GAN, and Diffusion. The semantic dimension is decreasing through filtering of principal components used to generate novel samples.
The VAEs conditioned on the same semantic representation exhibit a decline of 8% in density estimation for unseen classes in CUB as the dimensionality is reduced. The GAN also suffers from a notable decline in SUN with a 10% decline.
Figure 5: Conceptual understanding of cross-examining datasets. The semantic description of black and white stripes found on bird species is transferred from the bird dataset CUB to the animal dataset AWA.
Comparative analysis of performance across various domains. As for humans, classifying different species of birds requires finer scaled knowledge than classifying different species of animals. By cross-examining the datasets, we show that our approach not only reduced the training impact within a domain but actually can be reused across granularity, further reducing the overall training required. Our experiments show that our model is able to capture a few labels well in different datasets while missing other labels. We found that classes with semantic overlaps can be transferred across datasets, see Fig. (5). However, where the overlap is of no semantic similarity, i.e., sceneries in SUN, the performance is significantly reduced. Table 3 reports the average accuracy across unseen classes. We see that the model also performs better when the dimensional space of the test dataset is smaller, i.e. for APY. We recognise that performance is significantly reduced but argue the model is still utilizing learned knowledge where applicable.
To the best of our knowledge, there has been limited cross-examination of datasets in ZSL research, providing little foundation for direct comparison. We argue that training across different source domains offers valuable insights and suggest that future work should explore the characteristics of classes that are more easily transferable.
Table 3: Average accuracy for unseen classification in the datasets. Blue shading denotes that the model was trained in a reduced-dimensional space to accommodate the testing data, while green shading indicates that the model was trained in the original dimensional space but tested on a reduced-dimensional dataset to align with the trained model.
% Trained on
Tested on CUB SUN AWA APY
CUB - 1.6 15.1 7.9
SUN 1.3 - 7.8 1.0
AWA 11.1 3.9 - 1.5
APY 8.0 7.9 10.2 -

### v2

@@ 3.4 Model design
Our denoising Diffusion model employs a U-Net architecture, as introduced by the probabilistic diffusion model in [19]. To merge visual and semantic information effectively, we have customized this architecture to support both our time-dependent and visual-dependent embeddings, as illustrated in Fig.  2(b). To our knowledge, this represents the first application of a U-Net architecture tailored for zero-shot learning in such a specific way. The encoder-decoder structure of our U-Net is composed of linear blocks featuring ReLU non-linearity and batch normalization. Inputs to each layer include sinusoidal time embeddings and conditional data, which are extracted using self-attention mechanisms and augmented by a skip-connection between the encoding and decoding stages.
@@
for $s,\mathbf{x}\sim p(x,s)\in\mathcal{D}^{seen}$ do
[…]
     $s_{0}\leftarrow 2s-1$
@@
$\nabla_{\theta}\ [\lambda_{1}\mathcal{L}_{Diff}(\hat{s}_{0})+\lambda_{2}\mathcal{L}_{cls}(\hat{y})+\lambda_{3}\mathcal{L}_{noise}(\hat{\epsilon}_{t})]$
@@ 3.5 Sampling
Using standard methods from diffusion theory [19], we generate the semantic embedding space of an unseen sample through iterative conditional denoising using our trained model, as shown in Algorithm ( 2). Samples are drawn from the standard normal prior $p(s_{T})\sim\mathcal{N}(0,\mathbf{I})$ and denoised conditioned on the sinusoidal time-step embedding $\bar{t}_{i}\ \forall i\in[1000,0]$, and the Transformer-encoded latent visual space $x\in\mathcal{D}^{unseen}$.
@@
3.5.1 Classifier-free guidance.
To improve semantic alignment during sampling, we adopt the classifier-free guidance framework [20]. Using the difference between conditional and unconditional diffusion kernel estimates, we effectively steer the reverse process toward more probable samples under a desired condition. Using Tweedie’s formula [42], the denoising model can approximate the conditional and unconditional score functions $s_{\theta}(s_{t},t,x)\approx\nabla_{s_{t}}\log p(s_{t}|x)$ $s_{\theta}(s_{t},t,\emptyset)\approx\nabla_{s_{t}}\log p(s_{t})$.
By leveraging the decomposition of the conditional score we can mirror a kind of gradient in the semantic space, pulling the diffusion kernel in a direction that increases likelihood under the condition.
$\nabla_{s_{t}}\log p(s_{t}|x)=\nabla_{s_{t}}\log p(s_{t})+\nabla_{s_{t}}\log p(x|s_{t}),$ (19)
where $\nabla_{s_{t}}\log p(x|s_{t})$ acts as a classifier-like guidance term and can be reinterpreted as the difference between the conditional and unconditional scores:
$\nabla_{s_{t}}\log p(x|s_{t})\approx s_{\theta}(s_{t},t,x)-s_{\theta}(s_{t},t,\emptyset).$ (20)
To control the influence of the condition $x$, we scale this difference by a factor $g\in\mathbb{R}^{+}$, leading to the guided score estimate:
$s_{\theta}(s_{t},t,x)=s_{\theta}(s_{t},t,\emptyset)+g\cdot\left[s_{\theta}(s_{t},t,x)-s_{\theta}(s_{t},t,\emptyset)\right].$ (21)
This can be rearranged as our reversed transition function.
$s_{\theta}(s_{t},t,x)=(1+g)\cdot s_{\theta}(s_{t},t,x)-g\cdot s_{\theta}(s_{t},t,\emptyset).$ (22)
[…]
@@
$s_{t}\sim\mathcal{N}(0,\mathbf{I}),\ x\sim p(x)\in\mathcal{D}^{unseen},\ g:\textit{guidance strength}$
@@
     $s_{t}^{c},s_{t}^{u}\leftarrow Unet(s_{t},\bar{t},x),\ Unet(s_{t},\bar{t},\emptyset)$
     $\hat{s}\leftarrow(1+g)s_{t}^{c}-gs_{t}^{u}$
@@ 3.6 Zero-Shot Inference
In the zero-shot learning setting, the model utilizes the denoising model $s_{\theta}$ to approximate the semantic distribution given the instance $x^{unseen}$. A pseudo sample drawn from this distribution is then classified using a nearest-neighbor approach in the semantic space, leveraging the semantic density to bridge the gap between visual features of $x^{unseen}$ and class labels $y^{unseen}$:
$\hat{y}=\arg\min_{y\in\mathcal{Y}_{unseen}}\text{dist}(\hat{s}^{unseen},s^{unseen}_{y}),$ (23)
where $\hat{y}$ is the predicted class label for an unseen class instance, and $\text{dist}(\cdot,\cdot)$ denotes the distance metric, in our case cosine similarity:
@@ 4 Experimental results
Our analysis of diffusion as a generative method for zero-shot inference employs four publicly available benchmark datasets, each widely used in the field. This allows us to make a fair comparison of the quality and semantic coverage of the approximations. The benchmark datasets used are: Animals with Attributes 2 dataset (AwA2) [51], Caltech Birds dataset (CUB200-2011) [45], and Scene Understanding Attribute dataset (SUN) [35].
@@
The CUB dataset, focused on bird species, offers detailed representations in both image and semantic spaces. The semantic space consists of detailed attribute descriptions of the bird species’ characteristics. In contrast, AwA, which covers a range of different animal species, provides much coarser semantic descriptions, representing higher-level visual features. SUN consists of a scenery dataset with a wide range of classes, and its semantic descriptions are based on word2vec representations of each scene.
@@
Visual features are derived using a ResNet101 backbone pre-trained on ImageNet [50]. We only compare models using similar image features to ensure a fair evaluation. We use the semantic attributes released with each dataset, which are derived from either crowd-sourced human annotations or word2vec-based label extractions.
4.0.2 Implementation details.
@@
The employed U-net architecture for our Diffusion model consists of three hidden, fully connected dense layers, ReLU activation functions, and dropout for regularization. We use a feature extractor with a multi-head self-attention layer (MSA) for the conditional space. In the encoder and the decoder of the U-net, we concatenate and add the sinusoidal time embedding to layer inputs and the conditional features as explained in section 3.3.1. For our loss function, we fix $\lambda_{1}=1$ and $\lambda_{2}=1$ during training, while $\lambda_{3}$ varies depending on the dataset (see section 4.3).
4.1 Generalized accuracy
Our method achieves strong performance in the generalized zero-shot learning (GZSL) setting, particularly excelling on seen (S) class classification across all datasets. See Table  1. On AWA, we achieve the highest seen accuracy of 94.5%, substantially outperforming the previous best. Similarly, for CUB and SUN, we report 87.5% and 66.9% on seen classes, again outperforming existing state-of-the-art methods. We attribute this to the conditional nature of our diffusion model. Unlike prior approaches that condition on semantics to generate visuals, our reversed formulation conditions on visual inputs, allowing the model to preserve fine-grained visual distinctions. This becomes especially beneficial in datasets like AWA and SUN, where visual cues are rich and diverse. This highlights the advantage of diffusion models in capturing complex visual semantics through iterative refinement.
Table 1: Generalized zero-shot learning (GZSL) results on AWA, CUB, and SUN datasets. S, U, and H denote accuracy on seen classes, unseen classes, and their harmonic mean, respectively. Our method achieves the highest seen class accuracy across all datasets (such as 94.5% on AWA, 87.5% on CUB), and a competitive harmonic mean, especially on SUN (52.6%), highlighting the effectiveness of our conditional diffusion formulation.
AWA CUB SUN
Model Venue S U H S U H S U H
f-VAEGAN-D2 CVPR(19) 70.6 57.6 63.5 60.1 48.4 53.6 38.0 45.1 41.3
LisGAN CVPR(19) 76.3 52.6 62.3 57.9 46.5 51.6 37.8 42.9 40.2
GDAN ICCV(19) 67.5 32.1 43.5 66.7 39.3 49.5 40.9 38.1 53.4
ZSML IAAA(20) 58.9 74.6 65.8 60.0 52.1 55.7 45.1 21.7 29.3
DAZLE CVPR(20) 75.7 60.3 67.1 59.6 56.7 58.1 24.3 52.3 33.2
SDGZSL ICCV(21) 73.6 64.6 68.8 66.4 59.9 63.0 - - -
DAA-ZSL 79.9 65.7 72.1 65.5 66.1 65.8 38.7 47.8 42.8
HSVA NeurIPS(21) 76.6 53.9 66.8 58.3 57.2 55.3 39.0 48.6 43.3
ICCE CVPR(22) 65.3 82.3 72.8 65.5 67.3 66.4 - - -
FREE+ESZSL ICLR(22) 78.0 51.3 61.8 60.4 51.6 55.7 36.5 48.2 41.5
TF-VAEGAN+ESZSL ICLR(22) 74.7 55.2 63.5 63.3 51.1 56.6 39.7 44.0 41.7
TransZero AAAI(22) 61.3 82.3 70.2 69.3 68.3 68.8 52.6 33.4 40.8
TDCSS CVPR(22) 59.2 74.9 66.1 44.2 62.8 51.9 - - -
ZLAP IJCAI(22) 76.3 74.7 75.5 32.4 25.5 28.5 48.1 47.2 47.7
SE-GZSL AAAI(23) 68.1 58.3 62.8 53.3 41.5 46.7 30.5 50.9 34.9
TPR NeurIPS(24) 87.1 76.8 81.6 41.2 26.8 32.5 50.4 45.4 47.8
MAIN WACV(24) 81.8 72.1 76.7 58.7 65.9 62.1 40.0 50.1 48.8
Ours 94.5 42.4 58.3 87.5 32.3 47.2 66.9 43.4 52.6
However, our approach slightly underperforms on unseen (U) classes compared to some state-of-the-art methods. For instance, on AWA, we report 42.4%, compared to 82.3% achieved by ICCE. This is a result of the semantic space being estimated from a single visual instance: since each visual sample may correspond to several valid semantic embeddings, the diffusion model may distribute probability mass across divergent regions of the semantic space. This inherent multi-modality poses challenges in datasets with coarse or ambiguous semantic descriptors (e.g., AWA). Notably, our method performs more competitively on SUN’s unseen classes (43.4%, close to the best-performing method at 50.9%), where the broad and distributed semantic space (based on word2vec scene embeddings) mitigates this effect. Overall, our method achieves competitive harmonic means across all datasets: 58.3% (AWA), 47.2% (CUB), and 52.6% (SUN), while showcasing the strength of diffusion-based conditional modeling for generalized classification tasks.
4.2 Posterior approximation
@@
Generative model Name AwA CUB SUN Seen Unseen Harm. Seen Unseen Harm. Seen Unseen Harm. VAE cVAE 72.6 54.4 62.2 59.9 47.0 52.7 - - - GAN GAN 82.4 24.7 38.1 44.4 31.3 36.8 43.3 29.0 31.4 Diffusion (ours) RevCD 94.5 42.4 58.3 87.5 32.3 47.2 66.9 43.4 52.6
These comparisons are summarized in Table 2. As shown, no single generative model consistently outperforms the others across all datasets when generating both seen and unseen samples. Notably, our method surpasses the other approaches in generating samples when measured by the harmonic mean. The most significant performance gap is observed in the semantically coarse SUN dataset, where our approach achieves a 20% improvement over GANs. A similar trend is evident in the class-diverse AWA dataset, albeit with smaller margins. In contrast, the CUB dataset, which features a wide variety of fine-grained semantic details, proves challenging for denoising approaches, making variational inference methods a more effective fit.
VAEs benefit from the tractable estimation of the posterior distribution, as evidenced by a 5.4% higher harmonic mean when both seen and unseen samples are drawn from tighter distributions, such as those observed in the CUB dataset, which emphasizes local descriptions. In contrast, GANs may underperform in this context, likely due to mode collapse in the posterior. Our Diffusion model performs moderately, achieving a 10.4% improvement over GANs.
Conversely, GANs implicitly learn the distribution through adversarial training, which encourages the generator to produce high-fidelity samples, as demonstrated in the AWA dataset, where attributes emphasize global image descriptions. However, the absence of explicit density estimation makes GANs susceptible to seen-unseen bias, leading to a preference for the seen distribution during inference, as observed in the SUN datasets. Our Diffusion model shows strong performance in generating samples when the seen distribution is discriminative and low-dimensional, as in the AWA. However, it struggles to maintain a tight lower bound on the true data distribution as the dimensionality of the semantic space increases, as evidenced by its performance in the CUB dataset.
@@
4.3 Effect of classification loss
We investigate the influence of the classification loss weight $\lambda_{3}$ on the performance of our diffusion-based model (see Fig. (4)). The classification term guides the forward process during kernel estimation by biasing the diffusion trajectory toward semantically dense regions. Specifically, it introduces a gradient-based constraint that guides the process toward class-consistent subspaces, refining the learned transition kernel to denoise representations aligned with class identity.
Our reversed process estimates the semantic density of each class. The effect of this weight is notably different across datasets. On AWA, where semantic representations are coarse and broadly defined, increased weighting of the classification loss substantially boosts seen class accuracy—reaching near-perfect performance across all values of $\lambda_{3}$. However, this comes at the expense of unseen accuracy, which decreases as the model overfits to seen class modes, likely due to the reduced specificity in the semantic space, Fig. (4a). In contrast, CUB, characterized by fine-grained and richly structured semantic attributes, maintains high seen accuracy even at moderate classification weights, with a more stable and gradual decline in unseen performance, Fig. (4b). SUN, which features a broader class diversity and distributed word2vec-based semantics, displays a well-balanced trade-off. Seen accuracy is maximized at lower $\lambda_{3}$ values, while unseen performance remains relatively stable before degrading at higher weights, Fig. (4c). The robustness observed in SUN can be attributed to the high semantic variability, which naturally regularizes the kernel estimation and prevents premature convergence toward over-specialized representations.
Figure 4: Seen and unseen accuracy as a function of the classification loss weight $\lambda_{3}$ across AWA, CUB, and SUN. Increasing $\lambda_{3}$ consistently improves seen class accuracy but degrades unseen performance, with the trade-off varying by dataset. AWA exhibits sharp overfitting due to coarse semantics, while CUB remains more stable under stronger supervision. SUN shows a more balanced response, reflecting its broader semantic diversity.
These results highlight that the optimal weight of the classification-guided diffusion kernel is dataset-dependent. Fine-grained datasets like CUB benefit from moderate supervision, while coarsely defined or semantically diverse datasets (AWA and SUN) require careful calibration to avoid collapsing the semantic diversity necessary for zero-shot generalization.